# Marina del Rey CA 90292, USA

### v1

@@ Abstract
Hallucinations Generative AI Large Language Model.
Documentation: https://bitly.ws/34L6U
@@ 2 Related Work
Based on performance metrics, modern LLMs are considered definitively superior compared to an earlier generation of language models like BERT. They can handle complex prompts and instructions, but this complexity has also resulted in more sophisticated hallucination patterns [5]. Researchers are already attempting to define and categorize hallucinations in LLMs in a rigorous way, but efforts are ad-hoc and do not rely on the formal machinery that is a strength of the Semantic Web [13, 37]. Nevertheless, such works do offer insights into hallucination causes, types, and potential solutions, even though they do not always take into account the hallucinations published in non-academic sources.
@@ 3.5 Ontology Publication and Maintenance
HALO is published under the open CC-BY 4.0 license in two formats: (1) a machine-readable (OWL) file following ontology-publication best practices [4], with the files available both in a GitHub repository [^2] https://github.com/navapatn/halo-ontology and a Zenodo [^3] https://zenodo.org/records/10279463 repository; (2) a human-readable document that allows for more conceptual debate on how to represent hallucination categories and metadata. We documented HALO using Widoco [10], which generates a series of interlinked HTML pages offering a human-readable version of HALO, derived directly from its content. This documentation is made publicly accessible online through a persistent identifier, allowing us to fulfill FAIR’s Findability (specifically, F1, F2, and F4) requirements.
@@ 4.3 Metadata Module
5 Technical Specifications
This section summarizes HALO’s technical specifications augmenting those mentioned in Section 3, including interoperability, indexing and availability, logical correctness and validation, and documentation.
Interoperability: We imported 7 external classes into HALO, as one of our goals in designing HALO is to represent hallucinations that involve real-world entities and concepts, which are already present in vocabularies like FOAF. Note that the core implementation of HALO is OWL-based in order to achieve FAIR’s Interoperability (I1) principle.
Indexing and availability: We release the ontology publicly, under an open CC BY 4.0 International license, in Zenodo  [^5] https://zenodo.org/records/10279463. HALO can be browsed using any web browser, and using visualization tools such as WebVOWL.
@@ 6.1 Hallucination Dataset and Sources
In our effort to model and evaluate hallucination instances observed in actual LLMs using HALO, we compiled 40 prompts known to induce hallucinatory responses in LLMs from a diverse set of Web sources. The prompts span multiple document types, including Reddit posts, news articles from The New York Times, and reports from Opengov, among others. During this collection process, we also compiled information on applicable metadata, such as the specific LLM known to have produced the hallucination, date when we found the hallucination, Web source link, name, and type (e.g., blog post, newspaper article). We tabulate descriptive statistics on this collected data in Table 1.
@@
Number of Prompts
40
[How many games did the New Jersey Devils win in 2014]
Number of Sources
9
[Reddit]
Number of Metadata Collected
5
[Source, Source_link, Date, etc.]
Number of Document Types
4
[Post]
Date found
[April 1, 2023]
@@
ChatGPT (v3.5) [^7] https://chat.openai.com/ is a variant of the GPT (Generative Pretrained Transformer) family of generative language models developed by OpenAI, and has become widely known for its conversational abilities and its general knowledge.
BARD [^8] https://bard.google.com/chat is an LLM developed by Google AI. This model utilizes a transformer-based encoder-decoder architecture, incorporating a novel attention mechanism that allows it to effectively capture long-range dependencies within the input text.
Claude [^9] https://claude.ai/chats is an AI assistant developed by Anthropic. Similar to other LLMs, it was also designed to learn from human feedback and conversations, and to be more introspective about admitting its mistakes, refusing inappropriate requests, and discussing options that are more aligned with human ethics.
Because the original dataset we collected did not evaluate all prompts on all three LLMs, to showcase HALO’s utility, we first ‘completed’ the dataset by running all 40 prompts to all three LLMs, yielding a total of 120 data points. Figure 2 shows the knowledge graph of a data point from HALO. This served two purposes: first, it allowed us to verify that a hallucination was indeed an hallucination on the model on which it was documented; second, it allows us to pursue the stated objective through a competency query (CQ), as subsequently described. As shown in Table 2, even this relatively small dataset suggests that there is variation in the frequency and nature of hallucinations across the LLMs. A cursory analysis shows that, of the 40 prompts tested, BARD had the highest hallucination rate at 62.5%, followed by ChatGPT (42.5%) and Claude (40.0%). Particularly, BARD’s inability to refuse to respond to user prompts might have contributed to its higher rate of hallucinations. These verified hallucination instances, including metadata and corresponding responses from LLMs, form the foundational dataset for evaluating HALO and showcasing its utility using the CQs described in the next section.
@@
Statistic ChatGPT (GPT-3.5) BARD Claude
# Hallucinations
17
25
16
# Correct answers
13
15
13
# Refuse to answer
10
0
11
Hallucination Rate
42.5%
62.5%
40.0%
Figure 2: An illustrative example of a hallucination modeled by HALO
@@ 6.2 Ontology Evaluation
Id
Question Text
CQ1
What different types of generative AIs exhibit hallucination?
CQ2
What are the different types of hallucination that occurred by large language models :x?
CQ3
What is the ranking of hallucination types generated by model :x?
CQ4
Which large language model exhibits the most hallucination type :x?
CQ5
What pair of LLMs shared the same type of hallucinations the most?
@@
CQ
Input(s)
Expected Result(s)
CQ2
GPT-3.5
FF, FI LI, CI
CQ3
BARD
(1) FI, (2) FF, (3) CI, (4) II, (5) LI
CQ4
FactualityFabrication
BARD

### v2

# Further author information: (Send correspondence to M.K.)

@@ Abstract
Hallucination, Large Language Model, ChatGPT, Ontology
@@ 2 Related Work
Documented hallucinations in LLMs tend to be domain-specific, with question-answering (QA) being a prominent example of a task that can often elicit hallucinations. Numerous datasets and benchmarks have been developed for QA over the past year [23, 24, 25]. Additionally, efforts have been made to systematically generate hallucinations for ChatGPT in the QA domain, integrating with supervised human annotation to establish QA benchmarks [11]. The phenomenon of hallucination also extends to Large Vision-Language Models (LVLMs), where the model can hallucinate on producing an object and misidentify features that contradict textual and visual information [26]. These benchmarks are useful for task-specific prompts such as QA and text-visual generation, which motivated research in detecting and mitigating hallucinations, as illustrated in recent papers [20, 9, 27]. However, the interaction with newer or commercial LLMs like ChatGPT[28] often involves free-flowing text input, diverging from structured QA formats. This shift poses a challenge, as the free-form nature of interactions can lead to varied and unpredictable hallucination instances, making them harder to detect and document systematically. To address this, we are motivated in this paper to collect these instances and build a dataset and a structured framework, such as a hallucination ontology, to categorize and analyze these hallucinations systematically.
Based on performance metrics, modern LLMs are considered definitively superior compared to an earlier generation of language models like BERT. They can handle complex prompts and instructions, but this complexity has also resulted in more sophisticated hallucination patterns [29]. Researchers are already attempting to define and categorize hallucinations in LLMs in a rigorous way, but efforts are ad-hoc and do not rely on the formal machinery that is a strength of the Semantic Web [30, 12, 31]. Nevertheless, such works do offer insights into hallucination causes, types, and potential solutions, even though they do not always take into account the hallucinations published in non-academic sources.
@@ 3.5 Ontology Publication and Maintenance
HALO is published under the open CC-BY 4.0 license in two formats: (1) a machine-readable (OWL) file following ontology-publication best practices [40], with the files available both in a GitHub repository and a Zenodo repository (Section 5 provides the links and details); (2) a human-readable document that allows for more conceptual debate on how to represent hallucination categories and metadata. We documented HALO using Widoco[41], which generates a series of interlinked HTML pages offering a human-readable version of HALO, derived directly from its content. This documentation is made publicly accessible online through a persistent identifier, allowing us to fulfill FAIR’s Findability (specifically, F1, F2, and F4) requirements.
@@ 4.3 Metadata Module
5 Technical Specifications and Availability
We summarize HALO’s technical specifications augmenting those mentioned in Section 3, including interoperability, indexing and availability, logical correctness and validation, and documentation.
Interoperability: We imported seven external classes into HALO, as one of our goals in designing HALO is to represent hallucinations that involve real-world entities and concepts, which are already present in vocabularies like FOAF. Note that the core implementation of HALO is OWL-based in order to achieve FAIR’s Interoperability (I1) principle.
Indexing and availability: We release the ontology publicly [^3] https://bitly.ws/34Sdp, under a permissive CC BY 4.0 International License [^4] https://creativecommons.org/licenses/by/4.0/, as well as in Zenodo [^5] https://zenodo.org/records/10279463. HALO can be browsed using any web browser, and using visualization tools such as WebVOWL. It can also be accessed on GitHub [^6] https://github.com/navapatn/halo-ontology.
@@ 6.1 Hallucination Dataset and Sources
In our effort to model and evaluate hallucination instances, we contribute by constructing a compact dataset specifically designed to document hallucinatory responses. This dataset serves as an initial repository and facilitates the evaluation of HALO’s competency. We compiled 40 prompts known to induce hallucinatory responses in LLMs from a diverse set of Web sources. The prompts span multiple document types, including Reddit posts, news articles from The New York Times, and reports from Opengov, among others. During this collection process, we also compiled information on applicable metadata, such as the specific LLM known to have produced the hallucination, date when we found the hallucination, Web source link, name, and type (e.g., blog post, newspaper article). We tabulate descriptive statistics on this collected data in Table 1.
Figure 2: An illustration of the complete workflow for evaluating HALO. Hallucination instances are collected from multiple sources on the internet, typically as response examples from a single LLM. As described in Section 6.1, each prompt is executed across three different LLMs to complete the dataset, and instances are labeled as hallucinations when identified. Only the instances that exhibit hallucinations are then used to create Knowledge Graph (KG) instances, which are subsequently integrated into HALO for evaluation through SPARQL queries.
@@
Number of Prompts 40 [How many games did the New Jersey Devils win in 2014]
Number of Sources 9 [Reddit]
Number of Metadata Collected 5 [Source, Source_link, Date, etc.]
Number of Document Types 4 [Post]
Date found - [April 1, 2023]
@@
ChatGPT (v3.5)[43] is a variant of the GPT (Generative Pretrained Transformer) family of generative language models developed by OpenAI, and has become widely known for its conversational abilities and its general knowledge.
BARD[44] is an LLM developed by Google AI. This model utilizes a transformer-based encoder-decoder architecture, incorporating a novel attention mechanism that allows it to effectively capture long-range dependencies within the input text.
Claude[45] is an AI assistant developed by Anthropic. Similar to other LLMs, it was also designed to learn from human feedback and conversations, and to be more introspective about admitting its mistakes, refusing inappropriate requests, and discussing options that are more aligned with human ethics.
Figure 3: An illustrative example showing BARD’s responses that hallucinated when prompted in October 2023, but refused to answer in March 2024.
Because the original dataset we collected did not evaluate all prompts on all three LLMs, to showcase HALO’s utility, we first ‘completed’ the dataset by running all 40 prompts to all three LLMs. Figure 2 illustrates the complete process from dataset collection to HALO evaluation. We ran 40 prompts in both October 2023 and March 2024, not only to complete the dataset but also demonstrate that HALO can be used to document changes to hallucination instances over time. This resulted in 120 data points for each run, which yields a total of 240 data points. As shown in Table 2, the data from October 2023 suggests variations in both the frequency and nature of hallucinations across different LLMs, despite the dataset’s relatively small size. A cursory analysis shows that, of the 40 prompts tested, BARD had the highest hallucination rate at 62.5%, followed by ChatGPT (42.5%) and Claude (40.0%). Particularly, BARD’s inability to refuse to respond to user prompts might have contributed to its higher rate of hallucinations. In March 2024, the results for GPT-3.5 and Claude remained consistent with results from October 2023, while BARD exhibited a slight improvement, notably in refusing to answer questions on which it has no information. Figure 3 shows prompt examples where BARD used to hallucinate in October 2023, but improved its answer in later responses.
After completing the dataset, we generated knowledge graph (KG) instances for entries exhibiting hallucinations. Each KG instance includes all previously mentioned attributes and metadata, enhanced with hallucination labels corresponding to the category of hallucination identified in the responses. Figure 4 illustrates a KG representation of a data point from HALO. This served two purposes: first, it allowed us to verify that a hallucination was indeed a hallucination on the model on which it was documented; second, it allowed us to pursue the stated objective through a competency query (CQ), as subsequently described.
@@
March 2024 October 2023
Statistic GPT-3.5 BARD Claude GPT-3.5 BARD Claude
# Hallucinations 17 25 16 16 22 16
# Correct answers 13 15 13 14 15 13
# Refuse to answer 10 0 11 10 3 11
Hallucination Rate 42.5% 62.5% 40.0% 40.0% 55.0% 40.0%
Figure 4: An illustrative example of a hallucination modeled by HALO (in KG manner) using one data point from the dataset.
@@ 6.2 Ontology Evaluation
Id Question Text
CQ1 What different types of generative AIs exhibit hallucination?
CQ2 What are the different types of hallucination that occurred by large language models :x?
CQ3 What is the ranking of hallucination types generated by model :x?
CQ4 Which large language model exhibits the most hallucination type :x?
CQ5 What pair of LLMs shared the same type of hallucinations the most?
@@
CQ Input(s) Expected Result(s)
CQ2 GPT-3.5 FF, FI LI, CI
CQ3 BARD (1) FI, (2) FF, (3) CI, (4) II, (5) LI
CQ4 FactualityFabrication BARD